OpenStudy (anonymous):

I need formula Triangle ABC is similar to triangle DEF and has a ratio of similarity of 5:2. If side AB measures 30 mm, how long is side DE?again just the formula =)

jimthompson5910 (jim_thompson5910):

Since the ration of similarity is 5:2, this means that dividing any segment of ABC by the corresponding segment on DEF gives you 5/2 So AB/DE = 5/2

OpenStudy (anonymous):

i did 5/2 x 30/? giving me twelve is this ok too

jimthompson5910 (jim_thompson5910):

Well since AB/DE = 5/2 and AB = 30, this means 30/x = 5/2 where x is the length of DE

jimthompson5910 (jim_thompson5910):

if you solve that equation for x, you'll get x = 12, so you are correct

OpenStudy (anonymous):

ok good